# GoalRank: Group-Relative Optimization for a Large Ranking Model

### v1

@@ 1 Introduction
Recommender systems are indispensable for coping with the exponential growth of online content (Gomez-Uribe & Hunt, 2015). Industrial platforms typically adopt a multi-stage pipeline, comprising retrieval (He et al., 2020; Zhang et al., 2024) and ranking (Yu et al., 2019; Liu et al., 2023; Zhang et al., 2025). The ranking stage is particularly critical, as it determines the final sequence of items shown to users and has a major impact on both user satisfaction and platform revenue.
Formally, the ranking task can be defined as an $N{\!\rightarrow\!}L$ list-generation problem: given $N$ candidates from the preceding stage, the model outputs an ordered list of length $L$. The search space is the set of length-$L$ permutations, $\mathrm{P}(N,L)=\tfrac{N!}{(N-L)!}$, which makes exhaustive enumeration intractable for large $N$. Early approaches adopt a one-stage single generator that directly produces recommendation lists by scoring items and arranging them greedily (Zhuang et al., 2018; Ai et al., 2018; Pei et al., 2019a; Gong et al., 2022; Liu et al., 2023), as illustrated in Figure 1(a). However, this greedy strategy only models the item interdependencies in the candidate set (of size $N$) but not in the output list (of size $L$), often resulting in suboptimal rankings.
@@ 2. $i$
2 Preliminary
Let $\mathcal{U}=\{u\}$ be the user set and $\mathcal{V}=\{v\}$ be the item set. For each user $u\in\mathcal{U}$, let $\mathcal{V}_{u}\subset\mathcal{V}$ denote the candidate item set, with $|\mathcal{V}_{u}|=N$. Given a list length $L<N$, the list generation space for user $u$ is given by:
$\mathcal{L}_{u}:=\bigl\{\,l=(v_{1},\dots,v_{L})\in\mathcal{V}_{u}^{L}\,\big|\,v_{i}\neq v_{j}\ (i\neq j)\,\bigr\},\quad|\mathcal{L}_{u}|=\mathrm{P}(N,L):=\frac{N!}{(N-L)!}.$
Single-Stage Ranking. (G-only) A single-stage ranker directly generates a list through a generator. Let $\mathcal{X}_{u}$ denote the context information of user $u$ (e.g., historical interactions, side features). Given a ranking model $G(\cdot)$, the single-stage ranking can be formulated as:
@@ 2 Preliminary
Two-Stage Ranking. ((M-)G-E) In contrast, a two-stage ranking model employs an evaluator $E(\cdot)$ to assess the quality of candidate lists, and then selects the best one. Formally, the task can be expressed as
$l_{u}^{*}=\arg\max_{l\in\mathcal{L}_{u,k}}E(\mathcal{X}_{u},l),\quad\text{s.t.}\quad\mathcal{L}_{u,k}=\{\,l_{i}\mid l_{i}=G_{i}(\mathcal{X}_{u},\mathcal{V}_{u}),\ i=1,\dots,k\,\}.$
The single-generator–evaluator case corresponds to $k=1$, while the multi-generator–evaluator paradigm uses $k>1$ to enlarge the candidate-list space before evaluation. Due to constraints on space, a detailed discussion of related works is provided in Appendix A.
@@ 3.1 Can the generator-only paradigm perform better?
Proofs and technical details are deferred to Appendix B. Theorem 1 shows that for any two-stage ranking mixing $k$ small generators, there exists a sufficiently large one-stage generator-only ranking model whose induced policy space achieves a strictly smaller approximation error to $\pi^{*}$. Moreover, as the size of this generator increases (i.e., as $n$ grows), the approximation error can be driven arbitrarily close to zero. We remark that Theorem 1 is stated in terms of width scaling; the same conclusion holds under depth scaling, with proofs provided in Appendix B.
@@
Thus,
$\tau\log Z=\sup_{\pi}\Bigl\{\mathbb{E}_{l\sim\pi}\!\left[r^{*}(l)\right]+\tau\mathcal{H}(\pi)\Bigr\},$
@@
we can exploit this approximate order-invariance to construct a reference policy in a group-relative manner:
@@
where $\bar{r}_{\mathcal{B}}$ and $\sigma_{\mathcal{B}}$ are the mean and standard deviation of $\hat{r}$ over $\mathcal{B}$.
We then train a parametric policy $\pi_{\theta}$ to align with $\pi^{\mathrm{ref}}$ by minimizing the cross-entropy (equivalently, the KL divergence up to a constant): Finally, the training objective can be expressed as
@@ 3.3 GoalRank
Following Zhang et al. (2025), we first train a reward model $\hat{r}$ using real user feedback data, which can estimate the expected feedback for a given recommendation list (details are provided in Appendix C).
@@
As discussed in the previous section, constructing effective groups requires sufficiently large reward gaps among lists within each group, which is difficult to achieve when sampling multiple lists from a single generator. To address this, we introduce an auxiliary set of ranking policies $\mathcal{M}$ (including heuristic methods and lightweight neural models with implementation details provided in Appendix D). For each user $u$, we then construct a group of recommendation lists as
@@ 4.1.1 Datasets and Offline Experiments Setting
We conduct offline experiments on two public datasets, ML-1M (Harper, 2015) and Amazon-Book (McAuley et al., 2015), as well as two datasets of different scales collected from our industrial short-video platform, denoted as Industry and Industry-0.1B. The statistics of the four preprocessed datasets are summarized in Appendix E.1.
@@
Following common practices, we report Hit Ratio@L (H@L), NDCG@L (N@L), MAP@L (M@L), F1@L, and AUC with $L=6$. All reported results are averaged over five independent runs.
@@ 4.1.2 Baselines
We compare GoalRank against representative state-of-the-art methods from three major categories of ranking models:
  • Generator-only methods: These approaches rely on a single generator to produce item scores and directly generate the ranking list. Simple item-wise scoring models such as DNN (Covington et al., 2016) estimate user feedback independently for each user–item pair. More advanced methods, including DLCM (Ai et al., 2018), PRS (Feng et al., 2021a), PRM (Pei et al., 2019a), and MIR (Xi et al., 2022), explicitly capture mutual dependencies among candidate items.
@@
  • Multi-Generator–Evaluator methods: These approaches ensemble multiple generators to expand the candidate-list space and enhance ranking performance (Yang et al., 2025). In our experiments, we evaluate this strategy under different ensemble sizes, with the number of generators set to 3, 20, and 100.
To ensure fairness, all baselines are tuned within their respective parameter spaces. Unless otherwise specified (e.g., in scaling law experiments), the hidden embedding dimension of all models is fixed at 128, and model depths are kept consistent. Further details are provided in Appendix E.2.
@@ 4.1.3 Main Results (RQ1)
Improv. +17.12% +9.79% +15.43% +16.60% +7.98% +25.39% +20.15% +29.63% +28.04% +7.85% +4.07% +3.32% +5.60% +4.18% +2.19%
Ranking Performance. Table 1 reports the overall results across three datasets. We highlight three key observations:
  • GoalRank consistently achieves the best performance. Across all datasets and metrics (H@6, N@6, M@6, F1@6, and AUC), GoalRank outperforms all baselines. On ML-1M, it improves H@6 and M@6 by +17.12% and +15.43%; on the large Industry dataset, the gains are even more pronounced, reaching +25.39% in H@6 and +29.63% in M@6. These results confirm that a single large generator can better capture global ranking signals than multi-stage alternatives.
@@
In short, while G-E and MG-E offer incremental gains, GoalRank’s generator-only design provides a more scalable and fundamentally stronger solution.
@@
Scaling Performance. In Section 3.1, we showed theoretically that GoalRank admits scaling laws. Here, we empirically validate this by varying hidden dimensions, layer depth, and attention heads, adjusting model size from 1M to 0.1B. We compare GoalRank with three representative baselines: DNN, PIER, and MG-E. For fairness, DNN and PIER are scaled in the same manner as GoalRank, while the size of MG-E is increased by enlarging the number of generators.
Figure 3 presents the scaling performance on the Industry-0.1B dataset with respect to Hit Ratio@6 (H@6), NDCG@6 (N@6), MAP@6 (M@6), and F1@6. We exclude AUC since GoalRank already achieves values above 0.98 even at small model sizes, though we observe further improvements as the model size increases, approaching 1.0. The model size ranges from 1M to 0.1B parameters. [^2] For very small models, training on the full dataset leads to unstable convergence. To ensure fair comparison, we proportionally sample the dataset for all models (including GoalRank) at the same parameter scale. We summarize the key findings as follows:
@@
  • DNN and PIER show weak scaling. Larger sizes yield only modest improvements, reflecting the inherent limitation of pointwise scoring in approximating the optimal policy.
@@ 4.2.1 Experimental Setup
We evaluate GoalRank through a large-scale online A/B test on a real-world short-video recommendation platform. The platform serves over half a billion daily active users, with an item pool of tens of millions of videos. The system follows a two-stage workflow: (i) retrieval, which selects $N$ candidate items from millions, and (ii) ranking, which generates a final recommendation list of length $L$. GoalRank is deployed in the ranking stage, where $N=120$ candidates are provided and $L=6$ items are exposed to each user. The overall workflow is illustrated in Figure 4 (Appendix).

### v2

@@ 1 Introduction
Recommender systems are indispensable for coping with the exponential growth of online content (Gomez-Uribe and Hunt, 2015). Industrial platforms typically adopt a multi-stage pipeline, comprising retrieval (He et al., 2020; Zhang et al., 2024) and ranking (Yu et al., 2019; Liu et al., 2023; Zhang et al., 2025). The ranking stage is particularly critical, as it determines the final sequence of items shown to users and has a major impact on both user satisfaction and platform revenue.
Formally, the ranking task can be defined as an $N{\!\rightarrow\!}L$ list-generation problem: given $N$ candidates from the preceding stage, the model outputs an ordered list of length $L$. The search space is the set of length-$L$ permutations, $\mathrm{P}(N,L)=\tfrac{N!}{(N-L)!}$, which makes exhaustive enumeration intractable for large $N$. Early approaches adopt a one-stage single generator that directly produces recommendation lists by scoring items and arranging them greedily (Zhuang et al., 2018; Ai et al., 2018; Pei et al., 2019b; Gong et al., 2022; Liu et al., 2023), as illustrated in Figure 1(a). However, this greedy strategy only models the item interdependencies in the candidate set (of size $N$) but not in the output list (of size $L$), often resulting in suboptimal rankings.
@@ 2. $i$
2 Related Work and Preliminaries
The ranking task in recommender systems can be formulated as an $N{\!\rightarrow\!}L$ list-generation problem: given $N$ candidate items, the model outputs an ordered list of length $L$. Let $\mathcal{U}$ and $\mathcal{V}$ denote the user and item sets. For each user $u$, the candidate set is $\mathcal{V}_{u}\subseteq\mathcal{V}$ with $|\mathcal{V}_{u}|=N$. The generation space is
$\mathcal{L}_{u}=\bigl\{(v_{1},\dots,v_{L})\in\mathcal{V}_{u}^{L}:v_{i}\neq v_{j}\ (i\neq j)\bigr\},\qquad|\mathcal{L}_{u}|=\mathrm{P}(N,L)=\tfrac{N!}{(N-L)!}.$
Since $\mathrm{P}(N,L)$ is exponentially large, existing works build approximate ranking paradigms to explore this space.
Single-Stage Generator-Only Models. A straightforward solution is to use a single generator $G$ that scores items and constructs a list greedily (Zhuang et al., 2018; Ai et al., 2018; Pei et al., 2019b; Gong et al., 2022; Liu et al., 2023; Feng et al., 2021b; Xi et al., 2022; Pei et al., 2019a). Classic models include DLCM (Ai et al., 2018) and PRM (Pei et al., 2019b), which refine item scores using local listwise features. Formally:
@@ 2 Related Work and Preliminaries
Although efficient, these methods typically underexplore inter-item dependencies and may not remain consistent with the conditioning of the original candidate set.
Two-Stage Generator–Evaluator Paradigm. To better address the combinatorial $N{\!\rightarrow\!}L$ search space, recent work adopts a two-stage (multi-)Generator–Evaluator (G–E) framework (Chen et al., 2022; Shi et al., 2023; Xi et al., 2024; Lin et al., 2024; Ren et al., 2024b; Wang et al., 2025b). A generator produces multiple candidate lists, and an evaluator scores them to select the best one:
$l_{u}^{*}=\arg\max_{l\in\mathcal{L}_{u,k}}E(\mathcal{X}_{u},l),\qquad\mathcal{L}_{u,k}=\{\,G_{i}(\mathcal{X}_{u},\mathcal{V}_{u})\mid i=1,\dots,k\,\}.$
The special case k=1k{=}1 reduces to a single-generator–evaluator model. Multi-generator extensions ($k>1$) aim to enlarge the proposal space (Yang et al., 2025), but empirical gains saturate rapidly as $k$ grows. This diminishing return suggests that merely increasing the number of generators is inefficient and highlights the need for fundamentally stronger listwise modeling.
Other Directions. Other concurrent efforts incorporate large language models (Ren et al., 2024a; Gao et al., 2024; Wu et al., 2024; Gao et al., 2025; Ren et al., 2025; Liu et al., 2025) or reinforcement learning (Feng et al., 2021c; Wang et al., 2024; 2025c; Wei et al., 2020). LLM-based methods leverage textual side information, while RL-based approaches decompose the listwise value function to align rankings with user utility.
@@ 3.1 Can the generator-only paradigm perform better?
Proofs and technical details are deferred to Appendix A. Theorem 1 shows that for any two-stage ranking mixing $k$ small generators, there exists a sufficiently large one-stage generator-only ranking model whose induced policy space achieves a strictly smaller approximation error to $\pi^{*}$. Moreover, as the size of this generator increases (i.e., as $n$ grows), the approximation error can be driven arbitrarily close to zero. We remark that Theorem 1 is stated in terms of width scaling; the same conclusion holds under depth scaling, with proofs provided in Appendix A.
@@
$\text{Thus,}\qquad\tau$ $\log Z=\sup_{\pi}\Bigl\{\mathbb{E}_{l\sim\pi}\!\left[r^{*}(l)\right]+\tau\mathcal{H}(\pi)\Bigr\},$
@@
we can exploit this order-invariance to construct a reference policy in a group-relative manner:
@@
where $\bar{r}_{\mathcal{B}}$ and $\sigma_{\mathcal{B}}$ are the mean and standard deviation of $\hat{r}$ over $\mathcal{B}$. We then train a parametric policy $\pi_{\theta}$ to align with $\pi^{\mathrm{ref}}$ by minimizing the cross-entropy (equivalently, the KL divergence up to a constant): Finally, the training objective can be expressed as
@@ 3.3 GoalRank
Following Zhang et al. (2025), we first train a reward model $\hat{r}$ using real user feedback data, which can estimate the expected feedback for a given recommendation list (details are provided in Appendix B).
@@
As discussed in the previous section, constructing effective groups requires sufficiently large reward gaps among lists within each group, which is difficult to achieve when sampling multiple lists from a single generator. To address this, we introduce an auxiliary set of ranking policies $\mathcal{M}$ (including heuristic methods and lightweight neural models with implementation details provided in Appendix C). For each user $u$, we then construct a group of recommendation lists as
@@ 4.1.1 Datasets and Offline Experiments Setting
We conduct offline experiments on two public datasets, ML-1M (Harper, 2015) and Amazon-Book (McAuley et al., 2015), as well as two datasets of different scales collected from our industrial short-video platform, denoted as Industry and Industry-0.1B. The statistics of the four preprocessed datasets are summarized in Appendix D.1.
@@
Following common practices, we report Hit Ratio@L (H@L), NDCG@L (N@L), MAP@L (M@L), F1@L, and AUC with $L=6$. Reported results are averaged over five independent runs.
@@ 4.1.2 Baselines
We compare GoalRank against representative state-of-the-art methods from:
  • Generator-only methods: These approaches rely on a single generator to produce item scores and directly generate the ranking list. Simple item-wise scoring models such as DNN (Covington et al., 2016) estimate user feedback independently for each user–item pair. More advanced methods, including DLCM (Ai et al., 2018), PRS (Feng et al., 2021a), PRM (Pei et al., 2019b), and MIR (Xi et al., 2022), explicitly capture mutual dependencies among candidate items. We additionally compare with RankMixer (Zhu et al., 2025).
@@
  • Multi-Generator–Evaluator methods: These approaches ensemble multiple generators to expand the candidate-list space and enhance ranking performance (Yang et al., 2025). We evaluate this strategy under different ensemble sizes, with the number of generators set to 3, 20, and 100.
To ensure fairness, all baselines are tuned within their respective parameter spaces. Unless otherwise specified (e.g., in scaling law experiments), the hidden embedding dimension of all models is fixed at 128, and model depths are kept consistent.Moreover, all baselines share exactly the same evaluator (reward model) as GoalRank. Additional details on baseline configurations and the architecture of GoalRank are provided in Appendix D.2.
@@ 4.1.3 Main Results (RQ1)
Improv. +17.12% +9.79% +15.43% +16.60% +5.59% +25.39% +20.15% +29.63% +28.04% +7.73% +4.07% +3.32% +5.60% +4.18% +2.19%
Ranking Performance. Table 1 reports the overall results across three datasets. We highlight:
  • GoalRank consistently achieves the best performance. GoalRank outperforms all baselines. On ML-1M, it improves H@6 and M@6 by +17.12% and +15.43%; on the Industry dataset, the gains are even more pronounced, reaching +25.39% in H@6 and +29.63% in M@6. These results confirm that a single large generator can better capture ranking signals than multi-stage model.
@@
Scaling Performance. In Section 3.1, we showed theoretically that GoalRank admits scaling laws. Here, we empirically validate this by varying hidden dimensions, layer depth, and attention heads, adjusting model size from 1M to 0.1B. We compare GoalRank with four representative baselines: DNN, RanMixer, PIER, and MG-E. For fairness, baselines are scaled in the same manner as GoalRank, while the size of MG-E is increased by enlarging the number of generators.
Figure 3 presents the scaling performance on the Industry-0.1B dataset. We exclude AUC since GoalRank already achieves values above 0.98 even at small model sizes, though we observe further improvements as the model size increases, approaching 1.0. The model size ranges from 1M to 0.1B parameters. [^2] For very small models, training on the full dataset leads to unstable convergence. To ensure fair comparison, we proportionally sample the dataset for all models (including GoalRank) at the same parameter scale. We summarize the key findings as follows:
@@
  • Baselines show weak scaling. Larger sizes yield only modest improvements, reflecting the inherent limitation of pointwise scoring in approximating the optimal policy.
@@ 4.2.1 Experimental Setup
We evaluate GoalRank through a large-scale online A/B test on a real-world short-video recommendation platform. The platform serves over half a billion daily active users, with an item pool of tens of millions of videos. The system follows a two-stage workflow: (i) retrieval, which selects $N$ candidate items from millions, and (ii) ranking, which generates a final recommendation list of length $L$. GoalRank is deployed in the ranking stage, where $N=120$ candidates are provided and $L=6$ items are exposed to each user. The workflow and latency is illustrated in Figure 4 (Appendix).